OpenStudy (anonymous):

Integral Question: Int(tan(3x)^2*sec(3x)^2 dx please show full steps

OpenStudy (bahrom7893):

use wolframalpha it can show steps

OpenStudy (anonymous):

theres some parts of wolfram alpha that do it in a weird way

OpenStudy (bahrom7893):

which step do u not understand there?

OpenStudy (anonymous):

\[u = \tan(3x) \Rightarrow du = 3\sec^2(3x)dx\] \[\Rightarrow I = \int \tan^2(3x)\sec^2x(3x) = \frac{1}{3}\int u^2 du = \frac{u^3}{9}+K = \frac{\tan^3(3x)}{9} + K.\]